# Protecting Human Activity Signatures in Compressed IEEE 802.11 CSI Feedback

### v1

@@ IV Numerical Simulations
Figure 6: Constellation comparison for a single spatial stream under (a) deterministic midpoint quantization of the Givens angles, and (b) the proposed differentially private stochastic quantization (DP–SQ) mechanism for $N_{t}=2$, $N_{r}=1$, $16$-QAM, $\operatorname{SNR}=15$ dB and $\varepsilon=0.1$, $B_{\phi}=B_{\psi}=1$ bit.
Figure 7: Relative beamforming gain for several IEEE 802.11 antenna configurations ($\times 1, 2×2$$\times 2$, $2\times 3$, $2\times 4$, $2\times 8$). The perfect SVD beamformer defines the reference gain $G=1$, where $\epsilon_{\phi}=\epsilon_{\psi}=0.1$, $B_{\phi}=6$ bits and $B_{\psi}=3$ bits. For the BeamDancer scheme [cominelli2024physical], we adopt the randomization parameters that ensure the adversary’s activity classification accuracy remains strictly below $50\%$.
@@ V Conclusion & Future Work
In this paper, we have presented a standards-compatible framework for private CSI feedback based on DP quantization of the Givens rotation and phase angles used in IEEE 802.11ac/ax compressed beamforming. By replacing deterministic angle quantization with an $\varepsilon$-DP stochastic mechanism, the proposed design preserves the existing feedback format while providing formal privacy guarantees against activity, identity, and environment inference attacks. Numerical simulations under realistic 802.11 configurations demonstrate that DP angle quantization significantly suppresses adversarial classification accuracy while incurring only modest beamforming loss at the access point. These results highlight that meaningful privacy can be delivered within the current 802.11 compressed feedback architecture without modifying frame formats, pilot structures, or channel sounding procedures. The same solution can in principle be also applied to other 802.11 standards that support higher frequency mmWave bands (such as 802.11ad/11ay) or standards that incorporate more advanced hardware. Thus, our approach can be readily adapted to future wireless systems. Thus, future work includes extending the mechanism to multi-user beamforming, hybrid-precoding architectures, RIS-assisted systems, and adaptive DP budget allocation across streaming feedback packets. Our findings open a new direction in integrating rigorous privacy mechanisms into commodity Wi-Fi standards to protect users from passive CSI-based sensing attacks.

### v2

@@ IV Numerical Simulations
Figure 8: Constellation comparison for a single spatial stream under (a) deterministic midpoint quantization of the Givens angles, and (b) the proposed differentially private stochastic quantization (DP-SQ) mechanism for $N_{t}=2$, $N_{r}=1$, $16$-QAM, $\operatorname{SNR}=15$ dB and $\varepsilon=0.1$, $B_{\phi}=B_{\psi}=1$ bit.
Figure 9: Relative beamforming gain for several IEEE 802.11 antenna configurations ($\times 1, 2×2$$\times 2$, $2\times 3$, $2\times 4$, $2\times 8$). The perfect SVD beamformer defines the reference gain $G=1$, where $\epsilon_{\phi}=\epsilon_{\psi}=0.1$, $B_{\phi}=6$ bits and $B_{\psi}=3$ bits. For the BeamDancer scheme [2], we adopt the randomization parameters that ensure the adversary’s activity classification accuracy remains strictly below $50\%$.
TABLE III: Speed-based activity classification zones used for the micro-Doppler privacy attack.
Zone Speed Range (m/s) Activity Label
1 $0\leq|v|<0.5$ Stationary
2 $0.5\leq|v|<2.5$ Walking
3 $2.5\leq|v|<5.0$ Jogging
4 $5.0\leq|v|<\infty$ Running
@@
(a) CSI from Deterministic Quantization
(b) CSI from DP-SQ
(c) CSI from DP–GSQ
Figure 12: Beamformed CSI amplitude across subcarrier index and time for a user transitioning through stationary, walking, jogging, and running activity zones.
Figure 13: Monte Carlo evaluation of the DP-SQ mechanism over N=1, [^0] N=1{,}000 independent trials with random speed profiles. The dashed red line indicates the deterministic baseline ($p=0$). Error bars denote $\pm 1\sigma$ across trials. (a) Average beamforming gain versus randomization probability $p$. (b) Median beamforming gain versus randomization probability $p$. (c) Average speed classification error versus randomization probability $p$.
(a) CSI from Deterministic Quantization
(b) CSI from DP–SQ
Figure 14: Beamformed CSI amplitude across subcarrier index and time for a user transitioning through stationary, walking, jogging, and running activity zones.
Impact on Packet Loss. Fig. 14a shows the packet loss rate as a function of the randomization probability $p$ for beamforming various gain thresholds, $G_{\min}$. All curves originate near zero at $p=0$ and increase monotonically, confirming that the communication cost of DP-SQ scales smoothly with the privacy level. At the most stringent threshold ($G_{\min}=0.9$), the loss rate rises steeply, reaching approximately $80\%$ at $p=0.7$ and exceeding $90\%$ at $p=1.0$. Relaxing the threshold to $G_{\min}=0.8$ reduces the loss to roughly $70\%$ at full randomization, while the $G_{\min}=0.5$ curve remains below $40\%$ even at $p=1.0$, indicating that links operating with comfortable SNR margins can tolerate aggressive randomization. Fig. 14b presents the time-resolved packet loss for the $G_{\min}=0.8$ threshold as the randomization probability increases. Notably, the loss events do not cluster into prolonged outages but instead appear uniformly distributed in time, suggesting that DP-SQ degrades individual snapshot gains rather than inducing correlated fading bursts.
@@ V Conclusion & Future Work
In this paper, we have presented a standards-compatible framework for private CSI feedback based on DP quantization of the Givens rotation and phase angles used in IEEE 802.11ac/ax compressed beamforming. By replacing deterministic angle quantization with an $\varepsilon$-DP stochastic mechanism, the proposed design preserves the existing feedback format while providing formal privacy guarantees against activity, identity, and environment inference attacks. Numerical simulations under realistic 802.11 configurations demonstrate that DP angle quantization significantly suppresses adversarial classification accuracy while incurring only modest beamforming loss at the access point. These results highlight that meaningful privacy can be delivered within the current 802.11 compressed feedback architecture without modifying frame formats, pilot structures, or channel sounding procedures. The same solution can in principle be also applied to other 802.11 standards that support higher frequency mmWave bands (such as 802.11ad/11ay) or standards that incorporate more advanced hardware. Thus, our approach can be readily adapted to future wireless systems. Thus, future work includes extending the mechanism to multi-user beamforming [8], hybrid-precoding architectures, RIS-assisted systems [7], and adaptive DP budget allocation across streaming feedback packets. Our findings open a new direction in integrating rigorous privacy mechanisms into commodity Wi-Fi standards to protect users from passive CSI-based sensing attacks.